OpenStudy (anonymous):

Find the standard form of the equation of the ellipse with the following characteristics. Foci: (+-3,0) Major Axis of Length: 18

OpenStudy (amistre64):

since the foci are along the x axis; we get: x^2 y^2 --- + --- = 1 ; so far :) r^2 r^2

OpenStudy (amistre64):

i should rewrite that like this: x^2 y^2 --- + --- = 1 a^2 b^2

OpenStudy (amistre64):

a^2 - b^2 = c^2 ; where c = the distance from the center a foci; and a = the distance from the center to a vertex...in this case its 9 i believe

OpenStudy (amistre64):

9^2 - 3^2 = b^2 81 - 8 = b^2 73 = b^2 81 = a^2 x^2 y^2 --- + --- = 1 i believe is the form 81 73

OpenStudy (anonymous):

oh ok I see how you got that! Except it would be 72 not 73. Thanks

OpenStudy (anonymous):

I was close the first time I tried it.

OpenStudy (amistre64):

lol.....3^2 doesnt = 8? since when?

OpenStudy (anonymous):

lol it's al good I got the idea lol